A silver coated sphere of radius 5 cm carries a total charge of 12 nC uniformly
distributed on its surface in free space. Calculate (a) |D| on the surface of the
sphere. (b) D external to the sphere and (c) total energy stored in the field.
1
Expert's answer
2020-10-07T07:33:19-0400

(a) According to Gauss's law:


"\int D\cdot d S=q_{en},\\\space\\
D=\frac{q_{en}}{4\pi r^2}=3.82\cdot10^{-7}\text{ C/m}^2."

(b) D external to the sphere can be found for a>r:

"D=\frac{q_{en}}{4\pi a^2}=9.55\cdot10^{-10}/a^2\text{ C/m}^2."

(c) The energy is


"E=\frac{1}{2}\int D\cdot EdV=\frac{1}{2}\int\epsilon_0E^2dV=\\\space\\
=13\space\mu\text{J}."